# Horizon Imagination: Efficient On-Policy Training in Diffusion World Models

### v1

@@ 3 Preliminaries
Consider a Partially Observable Markov Decision Process (POMDP) environment. Here, we consider a practical state-agnostic formulation, as the agent has no knowledge about the POMDP hidden state space. At each step $t=1,2,\ldots$, starting from some initial observation ${\mathbf{o}}_{1}$, the agent picks an action ${\mathbf{a}}_{t}$ and the environment evolves according to ${\mathbf{o}}_{t+1},r_{t},d_{t}\sim p({\mathbf{o}}_{t+1},r_{t},d_{t}|{\mathbf{o}}_{\leq t},{\mathbf{a}}_{\leq t})$, where $r_{t}\in\mathbb{R},d_{t}\in\{0,1\}$ are the reward and termination signals, respectively. This process repeats until a non-zero termination signal is observed. The agent’s goal is to maximize its expected return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}]$.
@@ 4 Method
Multi-step trajectory generation naturally alternates between producing an observation ${\mathbf{o}}_{t}$ and sampling an action ${\mathbf{a}}_{t}\sim\pi({\mathbf{a}}_{t}\mid{\mathbf{o}}_{\leq t},{\mathbf{a}}_{<t})$, as each step depends on the output of the previous one. However, since the generative process of diffusion methods involves multiple (costly) forward passes for generating each observation, the above sequential generation of a trajectory becomes highly sequential and expensive, both in time and in compute.
@@ 4.2 Horizon Imagination (World Model Inference)
Consider a set of $N$ discrete actions $\mathcal{A}=[N]$. At denoising time $\tau$, the policy induces a categorical distribution $\pi^{\tau}=\pi({\mathbf{a}}_{t}^{\tau}|{\mathbf{z}}_{\leq t}^{\tau},{\mathbf{a}}_{<t})\in\Delta^{N-1}$ where $\Delta^{N-1}$ is the standard (N1)(N-1)-simplex. Inspired by inverse transform sampling, our method leverages a single multivariate uniform sample that is consistently mapped to an action under the evolving distributions $\pi^{\tau}$.
At the beginning of the denoising process, we draw two random objects: (i) a vector $\omega\sim\mathcal{U}([0,1))^{N-1}$, and (ii) a random permutation $\rho:[N]\rightarrow[N]$ that fixes an order over the actions. At each denoising time $\tau$, the action is given by ${\mathbf{a}}^{\tau}={\mathbf{a}}(\pi^{\tau},\omega)$, ensuring that the same uniform sample $\omega$ yields consistent decisions across the evolving distributions $\pi^{\tau}$. The action ${\mathbf{a}}(\pi,\omega)$ is determined by scanning $\omega$’s coordinates:
@@
The sampling scheme ${\mathbf{a}}(\cdot,\cdot)$ satisfies the following properties:
@@
    For any distribution ${\bm{p}}\in\Delta^{N-1}$, the random variable ${\mathbf{a}}({\bm{p}},\omega)$ with $\omega\sim\mathcal{U}([0,1))^{N-1}$ is distributed according to $\bm{p}$}, i.e., ${\mathbf{a}}({\bm{p}},\cdot)\sim{\bm{p}}$.
@@
    $A\;=\;\{\omega\ |\ {\mathbf{a}}({\bm{p}},\omega)\neq{\mathbf{a}}({\bm{q}},\omega)\}.$
@@ 4.3 Actor-Critic Training
$A_{k+t}\log\pi({\mathbf{a}}_{k+t}^{\tau_{t}}|{\mathbf{z}}_{\leq k}^{1},{\mathbf{z}}_{k+1}^{\tau_{1}},\ldots,{\mathbf{z}}_{k+t}^{\tau_{t}},{\mathbf{a}}_{<k+t},\tau)+\eta\mathcal{H}(\pi_{k+t}^{\tau_{t}}),$
@@
6:  Sample first action ${\mathbf{a}}_{{\textnormal{k}}}\sim\pi({\mathbf{a}}_{{\textnormal{k}}}|{\mathbf{z}}_{\leq{\textnormal{k}}},{\mathbf{a}}_{<{\textnormal{k}}})$
@@
12:   Compute action distributions $\pi_{{\textnormal{k}}+t}^{\tau}\leftarrow\pi({\mathbf{a}}_{{\textnormal{k}}+t}^{\tau}|{\mathbf{z}}_{\leq{\textnormal{k}}+t}^{\tau},{\mathbf{a}}_{<{\textnormal{k}}+t}^{\tau},\tau_{\leq t})$
13:   Generate actions ${\mathbf{a}}_{{\textnormal{k}}+t}^{\tau_{t}}={\mathbf{a}}\left(\pi_{{\textnormal{k}}+t}^{\tau},\omega_{t}\right)$ $\triangleright$ (Eq. 2)
@@ Appendix C Algorithm Pseudocode
19:   Sample action ${\mathbf{a}}_{t}\sim\pi({\mathbf{a}}_{t}|{\mathbf{z}}_{\leq t})$
@@
41:  Compute denoiser outputs $v_{\theta}({\mathbf{z}}^{\tau},\tau,{\mathbf{a}})$ (single forward pass)
@@ Appendix E Proof for Proposition 1
$\Pr({\mathbf{a}}({\bm{p}},\cdot)=i)$ $=\Pr(\omega_{i}<\alpha_{i}({\bm{p}}))\ \prod_{j=1}^{i-1}\Pr(\omega_{j}\geq\alpha_{j}({\bm{p}}))$
@@
$\frac{1}{2}\|{\bm{p}}-{\bm{q}}\|_{1}\leq\Pr(A)=\Pr({\mathbf{a}}({\bm{p}},\omega)\neq{\mathbf{a}}({\bm{q}},\omega))\quad\forall\omega\in[0,1)^{N-1}$
@@
$A\;=\;\{\ {\mathbf{a}}(p,\omega)\neq{\mathbf{a}}(q,\omega)\},\quad\omega\sim\mathcal{U}([0,1))^{N-1}.$
@@
Notice that by the definition of ${\mathbf{a}}(\cdot,\cdot)$, it holds that ${\mathbf{a}}({\bm{p}},\omega)={\mathbf{a}}({\bm{q}},\omega)$ for all $\omega\in\left[\bigcup_{i=1}^{N-1}L_{i}\right]\cup U$. Similarly, $\Pr(A)=\Pr(\omega\in\bigcup_{i=1}^{N-1}C_{i})$. Hence,
@@ Appendix F Potential Extension to Continuous Action Spaces
$\mu_{t},\sigma_{t}=\pi({\mathbf{a}}_{t}|{\mathbf{z}}_{\leq t},{\mathbf{a}}_{<t}).$
@@
${\mathbf{a}}(\pi,\omega)=\mu+\omega\,\sigma.$
This construction would minimize action changes when the policy distribution parameters $\mu,\sigma$ remain fixed, reduce action differences when they vary, and ensure that ${\mathbf{a}}(\pi,\omega)$ follows the policy’s distribution, as in the discrete case. Notably, such an extension may also avoid the computational overhead associated with classifier-guidance methods used in prior work.

### v2

# Horizon Imagination: Efficient On-Policy Rollout in Diffusion World Models

@@ 3 Preliminaries
Consider a Partially Observable Markov Decision Process (POMDP) environment. Here, we consider a practical state-agnostic formulation, as the agent has no knowledge about the POMDP hidden state space. At each step $t=1,2,\ldots$, starting from some initial observation ${\mathbf{o}}_{1}$, the agent picks an action ${\textnormal{a}}_{t}$ and the environment evolves according to ${\mathbf{o}}_{t+1},r_{t},d_{t}\sim p({\mathbf{o}}_{t+1},r_{t},d_{t}|{\mathbf{o}}_{\leq t},{\mathbf{a}}_{\leq t})$, where $r_{t}\in\mathbb{R},d_{t}\in\{0,1\}$ are the reward and termination signals, respectively. This process repeats until a non-zero termination signal is observed. The agent’s goal is to maximize its expected return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}]$.
@@ 4 Method
Multi-step trajectory generation naturally alternates between producing an observation ${\mathbf{o}}_{t}$ and sampling an action ${\textnormal{a}}_{t}\sim\pi(\cdot\mid{\mathbf{o}}_{\leq t},{\mathbf{a}}_{<t})$, as each step depends on the output of the previous one. However, since the generative process of diffusion methods involves multiple (costly) forward passes for generating each observation, the above sequential generation of a trajectory becomes highly sequential and expensive, both in time and in compute.
@@ 4.2 Horizon Imagination (World Model Inference)
Consider a set of $N$ discrete actions $\mathcal{A}=[N]$. At denoising time $\tau$, the policy induces a categorical distribution $\pi^{\tau}=\pi(\cdot|{\mathbf{z}}_{\leq t}^{\tau},{\mathbf{a}}_{<t})\in\Delta^{N-1}$ where $\Delta^{N-1}$ is the standard (N1)(N-1)-simplex. Inspired by inverse transform sampling, our method leverages a single multivariate uniform sample that is consistently mapped to an action under the evolving distributions $\pi^{\tau}$.
At the beginning of the denoising process, we draw two random objects: (i) a vector $\omega\sim\mathcal{U}([0,1))^{N-1}$, and (ii) a random permutation $\rho:[N]\rightarrow[N]$ that fixes an order over the actions. At each denoising time $\tau$, the action is given by ${\textnormal{a}}^{\tau}={\textnormal{a}}(\pi^{\tau},\omega)$, ensuring that the same uniform sample $\omega$ yields consistent decisions across the evolving distributions $\pi^{\tau}$. The action ${\textnormal{a}}(\pi,\omega)$ is determined by scanning $\omega$’s coordinates:
@@
The sampling scheme ${\textnormal{a}}(\cdot,\cdot)$ satisfies the following properties:
@@
    For any distribution ${\bm{p}}\in\Delta^{N-1}$, the random variable ${\textnormal{a}}({\bm{p}},\omega)$ with $\omega\sim\mathcal{U}([0,1))^{N-1}$ is distributed according to $\bm{p}$}, i.e., ${\textnormal{a}}({\bm{p}},\cdot)\sim{\bm{p}}$.
@@
    $A\;=\;\{\omega\ |\ {\textnormal{a}}({\bm{p}},\omega)\neq{\textnormal{a}}({\bm{q}},\omega)\}.$
@@ 4.3 Actor-Critic Training
$A_{k+t}\log\pi({\textnormal{a}}_{k+t}^{\tau_{t}}|{\mathbf{z}}_{\leq k}^{1},{\mathbf{z}}_{k+1}^{\tau_{1}},\ldots,{\mathbf{z}}_{k+t}^{\tau_{t}},{\mathbf{a}}_{<k+t},\tau)+\eta\mathcal{H}(\pi_{k+t}^{\tau_{t}}),$
@@
6:  Sample first action ${\textnormal{a}}_{{\textnormal{k}}}\sim\pi(\cdot\,|\,{\mathbf{z}}_{\leq{\textnormal{k}}},{\mathbf{a}}_{<{\textnormal{k}}})$
@@
12:   Compute action distributions $\pi_{{\textnormal{k}}+t}^{\tau}\leftarrow\pi(\cdot\,|\,{\mathbf{z}}_{\leq{\textnormal{k}}+t}^{\tau},{\mathbf{a}}_{<{\textnormal{k}}+t}^{\tau},\tau_{\leq t})$
13:   Generate actions ${\textnormal{a}}_{{\textnormal{k}}+t}^{\tau_{t}}={\textnormal{a}}\left(\pi_{{\textnormal{k}}+t}^{\tau},\omega_{t}\right)$ $\triangleright$ (Eq. 2)
@@ Appendix C Algorithm Pseudocode
19:   Sample action ${\textnormal{a}}_{t}\sim\pi({\textnormal{a}}_{t}|{\mathbf{z}}_{\leq t})$
@@
41:  Compute denoiser outputs $v_{\theta}({\mathbf{z}}^{\tau},\tau,{\textnormal{a}})$ (single forward pass)
@@ Appendix E Proof for Proposition 1
$\Pr({\textnormal{a}}({\bm{p}},\cdot)=i)$ $=\Pr(\omega_{i}<\alpha_{i}({\bm{p}}))\ \prod_{j=1}^{i-1}\Pr(\omega_{j}\geq\alpha_{j}({\bm{p}}))$
@@
$\frac{1}{2}\|{\bm{p}}-{\bm{q}}\|_{1}\leq\Pr(A)=\Pr({\textnormal{a}}({\bm{p}},\omega)\neq{\textnormal{a}}({\bm{q}},\omega))\quad\forall\omega\in[0,1)^{N-1}$
@@
$A\;=\;\{\ {\textnormal{a}}(p,\omega)\neq{\textnormal{a}}(q,\omega)\},\quad\omega\sim\mathcal{U}([0,1))^{N-1}.$
@@
Notice that by the definition of ${\textnormal{a}}(\cdot,\cdot)$, it holds that ${\textnormal{a}}({\bm{p}},\omega)={\textnormal{a}}({\bm{q}},\omega)$ for all $\omega\in\left[\bigcup_{i=1}^{N-1}L_{i}\right]\cup U$. Similarly, $\Pr(A)=\Pr(\omega\in\bigcup_{i=1}^{N-1}C_{i})$. Hence,
@@ Appendix F Potential Extension to Continuous Action Spaces
$\mu_{t},\sigma_{t}=\pi(\cdot|{\mathbf{z}}_{\leq t},{\mathbf{a}}_{<t}).$
@@
${\textnormal{a}}(\pi,\omega)=\mu+\omega\,\sigma.$
This construction would minimize action changes when the policy distribution parameters $\mu,\sigma$ remain fixed, reduce action differences when they vary, and ensure that ${\textnormal{a}}(\pi,\omega)$ follows the policy’s distribution, as in the discrete case. Notably, such an extension may also avoid the computational overhead associated with classifier-guidance methods used in prior work.